Computing in Astronomy: Practical work 18.

Make an IDL program that plots two parallel horizontal lines. The lines should not touch the coordinate axes. Write vertically between the lines ’ω is omega’. Give the picture a title Practical work 18, name x-axis with your own name and y-axis with the date when you did the work.

Make a ps-file of the picture. Add the picture and the IDL program listing in the report that is made with LaTeX. If asked the student should be able to show the tex-file to teacher.